From GlueXWiki
Jump to: navigation, search

April 19, 2012 FDC meeting


  1. New results from the oxygen studies (Lubomir)
  2. Production Construction Tracking (Dave)
    • New procedures for package refurbishment
    • First production cell with Apiezon and Viton O-rings
    • Production status
  3. Engineering update (Bill)
  4. Electronics update (Chris)
  5. Tests with first package at 126 (Beni)
    • Grounding
  6. Other


Participants: Bill, Dave, Mark, Simon, Beni, Eugene, and Lubomir.

New results from the oxygen studies

- Latest results linked above, starting from configuration #19: applying vacuum grease (Apiezon (L)) on two EPDM O-rings in the testing chamber resulted in a record low oxygen of 50ppm. Without Apiezon with the same configuration we had 405ppm. It means we had significant oxygen contribution from the bottom cathode surface against the top O-ring on the spacer, that was eliminated with the grease. It turned out, however, the grease reacted with the EPDM swelling it and making it inelastic, but at the same time sealing the surfaces. We repeated the same experiment with Viton (that doesn't react with Apiezon) and got 70ppm oxygen. Based on this we decided to use Viton with Apiezon for the production.

- Oxygen estimate for the full package: Based on the latest tests one has to count not the number of the O-rings but the "bad" surfaces. The Lexan, plastic surfaces and the Hysol-ed ones are considered good. That's why we have one bad surface (bottom of the cathode) in the testing package and in the production cell we have 4. Assuming 40ppm comes from the supply gas, one bad surface contributes 30ppm, so with the same flow of 220ccpm we will have 120ppm. Therefore at 80ccpm per cell or 480ccpm total for the package, we expect 330ppm (0.03%).


- Dave: Chris will put components on wire plane #24 tomorrow, #26 on the stringing table, three wire planes waiting for deadening. Working also on cathodes #45, 46 and 47. The end window with wavy Mylar was fixed: old foil removed and a new one tensioned.

- First production cell with Viton and Apiezon on all the O-rings was assembled yesterday. The oxygen goes down with relatively high rate and one day after the installation is ~3,000ppm. However, as before with Viton O-rings, we found significant gas leak. Again as before, we reduced it by two orders of magnitude using C-clamps, but Bill is against this since we can deform the gusset rings. After the meeting Bill looked at it at Blue Crab. Most likely the leak is because the gusset ring is not flat at that place. Bill will try to shim that place after we finish the oxygen tests, but for the future he proposed using hard rubber between the gusset ring and the end window.

- Dave is working on new procedures to install O-rings. In addition now we have to coat with Hysol the top surfaces of the G10 frames, then after installing the O-ring, we apply Apiezon on the top of the O-ring. Coating is needed in case we want to remove the Apiezon.


- With the grease on the O-rings package disassembling is more complicated. To avoid cleaning, changing all the O-rings, we discussed if we can remove several planes or a full cell together without the need to redo all the O-rings inside that section. Bill will work on such tooling. The idea is to pull the metal pins (one by one) down to the position where the section that we want to remove ends, and then to install other pins there only for that section.

- Bill presented the fiducial scheme to survey people. He will do some modifications: adding 4 new fiducial points on the gusset rings of the first and last package visible from outside. These is needed to make connection to the flags on the periphery of all the packages. Bill is working also on the package spacers.

- Bill did conceptual design of the helium drift chamber to be proposed on the GlueX upgrade workshop.


- Dave: Chris is busy with BCAL now; has not started working on the grounding waiting for Fernando to do noise measurements with Beni on the first package.

- Beni: We need to figure out with Fernando the LV supply that will be needed soon for the package tests.

Tests with first package at 126

- Beni set the threshold at 7.5V (before it was 13V) and tried to improve the grounding on the wire cards, looking at the noise with a scope one card at a time. He managed to reduce the noise on planes 6 to 3 now working on plane 2. He found it is important to connect the golden surfaces of the signal PCBs to the cathode ground below that wire plane. Grounding works with aluminum foil, but not with the green wires. Beni will try using the braid that Fernando suggested for grounding. If it works we can solder it to the golden surfaces. During the tests one pre-amp was burned, most likely the voltage regulator, by touching the end of the flat signal cable in the connector with aluminum foil.